OpenStudy (anonymous):

Suppose you know the graph of f contains the point (2,3). If g(x)=f(x+5). What point must be on the graph of g? Possible answers: (2,8) (-3,3) (7,2) (3,2) (2,-2)

OpenStudy (accessdenied):

Well, if you only know that f(2) = 3, then we can only define g for some x-value that gets us f(2) from f(x + 5).

OpenStudy (asnaseer):

as AccessDenied indicted, you know: f(2) = 3 so you need to get f(2) from f(x+5) which leads to solving the equation: x+5 = 2 this will give you a value for x such that: f(x+5) = f(2) = 3 you should then be able to deduce the value of g(x) at this point.